# LongRAG: Enhancing Retrieval-Augmented Generation with Long-context LLMs

### v1

@@ Abstract
In traditional RAG framework, the basic retrieval units are normally short. The common retrievers like DPR normally work with 100-word Wikipedia paragraphs. Such a design forces the retriever to search over a large corpus to find the ‘needle’ unit. In contrast, the readers only need to extract answers from the short retrieved units. Such an imbalanced ‘heavy’ retriever and ‘light’ reader design can lead to sub-optimal performance. In order to alleviate the imbalance, we propose a new framework LongRAG, consisting of a ‘long retriever’ and a ‘long reader’. LongRAG processes the entire Wikipedia into 4K-token units, which is 30x longer than before. By increasing the unit size, we significantly reduce the total units from 22M to 600K. This significantly lowers the burden of retriever, which leads to a remarkable retrieval score: answer recall@1=71% on NQ (previously 52%) and answer recall@2=72% (previously 47%) on HotpotQA (full-wiki). Then we feed the top-k retrieved units ($\approx$ 30K tokens) to an existing long-context LLM to perform zero-shot answer extraction. Without requiring any training, LongRAG achieves an EM of 62.7% on NQ and 64.3% on HotpotQA (full-wiki), which is on par with the SoTA model. Our study offers insights into the future roadmap for combining RAG with long-context LLMs.
@@ 1 Introduction
Figure 1: Traditional RAG vs. LongRAG. (Up) Traditional RAG operates on short retrieval units, where retriever needs to scan over massive amount of units to find the relevant piece. In contrast, LongRAG operates on long retrieval units (30x longer). Retriever has a much less workload, which significantly boosts the recall score. LongRAG fully exploits the ability of long-context language models (reader) to achieve strong performance. (Middle) LongRAG vs. other RAG methods on NQ. Blue model names indicate models without fine-tuning, while red model names indicate models with fine-tuning. (Down) LongRAG vs. other RAG methods on HotpotQA.
Retrieval-Augmented Generation (RAG) methods have long been employed to enhance large language models (LLMs) Mialon et al. (2023). Knowledge in the form of natural language can be entirely offloaded from the parametric knowledge of LLMs by leveraging a standalone retrieval component from an external corpus. The existing RAG framework tends to use short retrieval units, such as 100-word passages in popular open-domain question answering tasks Chen et al. (2017); Lewis et al. (2020); Karpukhin et al. (2020). The retriever is tasked with finding the “needle” (i.e. the precise tiny retrieval unit) from the “haystack” (i.e. the massive corpus with tens of millions of information units). Subsequently, the retrieved units are passed to the reader to generate the final response. On the contrary, the reader only needs to extract answers from these retrievals, which is a fairly easy task. This kind of imbalanced design, with a “heavy” retriever and a “light” reader, puts too much pressure on the retriever. Therefore, the state-of-the-art RAG models Izacard and Grave (2020b) need to recall huge amount of units, such as the top-100 or even more, combined with additional complex re-ranker to achieve great performance. Moreover, short retrieval units can lead to semantic incompleteness due to document truncation. This can lead to information loss, ultimately restricting the end performance. This traditional design choice of the RAG framework was made in an era when NLP models were heavily restricted by their ability to handle long contexts. With the recent advances in long-context language models, the reader can potentially handle up to 128K or even millions of tokens as input Reid et al. (2024); Achiam et al. (2023). In this paper, we propose to revisit this design choice for open-domain question answering and propose LongRAG framework as a solution to balance the workload between the retriever and the reader, as illustrated in Figure 1. There are three important designs in our novel framework:
@@
    Long Retrieval Unit: By using entire Wikipedia documents or grouping multiple related documents, we can construct long retrieval units with more than 4K tokens. This design could also significantly reduce the corpus size (number of retrieval units in the corpus). Then, the retriever’s task becomes much easier. Additionally, the long retrieval unit will also improve the information completeness to avoid ambiguity or confusion.
@@
    Long Retriever: The long retriever will identify coarse relevant information for the given query by searching through all the long retrieval units in the corpus. The top 4 to 8 retrieval units are concatenated as the retrieved long context for the next step.
@@
    Long Reader: The long reader will further extract answers from the concatenation of retrievals, which is normally around 30K tokens. We simply prompt an existing long-context LM (like Gemini or GPT4) with the question to produce the answers.
These three novel designs significantly boost the overall performance of RAG on open-domain question answering tasks like NQ Kwiatkowski et al. (2019) and HotpotQA Yang et al. (2018). LongRAG has several advantages: 1) It does not require additional re-rankers, and the best results can be attained by only considering the top 4-8 retrieved units. 2) The long retrieval unit amalgamates comprehensive information from related documents, which can be used directly to answer multi-hop questions without iterative retrieval.
In our experiments, we adopt off-the-shelf retriever like BGE Xiao et al. (2023) and reader like Gemini-1.5-Pro (Reid et al., 2024) or GPT-4o OpenAI (2024) without any tuning on NQ or HotpotQA. In our experiments, we reduce the NQ corpus size from 22M to 600K document units, which improves the answer recall@1 from 52% (DPR) to 71%. Similarly, we reduce the HotpotQA corpus size from 5M to 500K, which improves the recall@2 from 47% (DPR) to 72%. The improvement in retriever can significantly benefit the reader model. By exploiting the long-context understanding ability of GPT-4o, LongRAG can achieve an EM of 62% on NQ and 64% on HotpotQA. These results could be comparable to the strongest fine-tuned RAG models like Atlas Izacard et al. (2022) and MDR Xiong et al. (2020b). Through these experiments, we have verified the effectiveness of LongRAG in handling open domain question answering. We believe LongRAG could provide strong insights into how we build modern RAG system. Meanwhile, there is still room for improvement in our framework, particularly the need for stronger long embedding models, as shown in Table 3. Additionally, more general methods to formulate long retrieval units beyond hyperlinks will be helpful.
@@ 3.1 Long Retriever
We approximate it by maximizing the scores of all chunks $g^{\prime}$ within the retrieval unit $g$. We consider different levels of granularity, including passage level, document level, and the complete grouped document. The empirical study about this settings is in Table 3. With this similarity score setup, we will retrieve the top $k$ retrieval units closest to the given query. For efficient retrieval, we precompute the embedding of each retrieval unit $g^{\prime}$ and predict the exact inner product search index in FAISS (Johnson et al., 2019).
@@ 4.2 Retrieval Performance
BGE-Large 512-tokens passage 71.7%
E5-Mistral-7B document (Avg 4K tokens) 54.2%
E5-Mistral-7B grouped documents (Avg 6K tokens) 19.6%
@@ 4.3 Full QA Performance
In Figure 5, we compare the performance of three different readers: Gemini-1.5-pro, GPT-4-Turbo, and GPT-4o. The results indicate that GPT-4o achieves the highest exact match score on the 200 test questions of the NQ dataset among the three models. This suggests that GPT-4o is the most effective in the role of a long reader in the LongRAG framework. The enhanced performance of GPT-4o can be attributed to its superior ability to process and comprehend lengthy contexts, ensuring that crucial information is accurately extracted. Therefore, we mainly report the GPT-4o results in our main table.

### v2

@@ Abstract
In traditional RAG framework, the basic retrieval units are normally short. The common retrievers like DPR normally work with 100-word Wikipedia paragraphs. Such a design forces the retriever to search over a large corpus to find the ‘needle’ unit. In contrast, the readers only need to extract answers from the short retrieved units. Such an imbalanced ‘heavy’ retriever and ‘light’ reader design can lead to sub-optimal performance. In order to alleviate the imbalance, we propose a new framework LongRAG, consisting of a ‘long retriever’ and a ‘long reader’. LongRAG processes the entire Wikipedia into 4K-token units, which is 30x longer than before. By increasing the unit size, we significantly reduce the total units from 22M to 600K. This significantly lowers the burden of retriever, which leads to a remarkable retrieval score: answer recall@1=71% on NQ (previously 52%) and answer recall@2=72% (previously 47%) on HotpotQA (full-wiki). Then we feed the top-k retrieved units ($\approx$ 30K tokens) to an existing long-context LLM to perform zero-shot answer extraction. Without requiring any training, LongRAG achieves an EM of 62.7% on NQ and 64.3% on HotpotQA (full-wiki), which is on par with the (fully-trained) SoTA model. Our study offers insights into the future roadmap for combining RAG with long-context LLMs.
@@ 1 Introduction
Figure 1: Traditional RAG vs. LongRAG. (Up) Traditional RAG operates on short retrieval units, where the retriever needs to scan over a massive amount of units to find the relevant piece. In contrast, LongRAG operates on long retrieval units (30x longer). Retriever has a much less workload, which significantly boosts the recall score. LongRAG fully exploits the ability of long-context language models to achieve strong performance.
Retrieval-Augmented Generation (RAG) methods have long been employed to enhance large language models (LLMs) Mialon et al. (2023). Knowledge in the form of natural language can be entirely offloaded from the parametric knowledge of LLMs by leveraging a standalone retrieval component from an external corpus. The existing RAG framework tends to use short retrieval units, such as 100-word passages in popular open-domain question-answering tasks Chen et al. (2017); Lewis et al. (2020); Karpukhin et al. (2020). The retriever is tasked with finding the “needle” (i.e. the precise tiny retrieval unit) from the “haystack” (i.e. the massive corpus with tens of millions of information units). Subsequently, the retrieved units are passed to the reader to generate the final response. On the contrary, the reader only needs to extract answers from these retrievals, which is a fairly easy task. This kind of imbalanced design, with a “heavy” retriever and a “light” reader, puts too much pressure on the retriever. Therefore, existing RAG models Izacard and Grave (2020b) have to recall huge amounts of units, such as the top-100/200, combined with additional re-ranker to achieve the best performance. Moreover, short retrieval units can lead to semantic incompleteness due to document truncation. This can lead to information loss, ultimately hurting the end performance. This design choice was made in an era when NLP models were heavily restricted by their ability to handle long contexts. With the recent advances in long-context language models, the retriever and reader can potentially handle up to 128K or even millions of tokens as input Reid et al. (2024); Achiam et al. (2023).
In this paper, we propose to revisit this design choice for open-domain question answering and propose the LongRAG framework as a solution to balance the workload between the retriever and the reader, as illustrated in Figure 1. There are three important designs in our novel framework:
@@
    Long Retrieval Unit: By using entire Wikipedia documents or grouping multiple related documents, we can construct long retrieval units with more than 4K tokens. This design could also significantly reduce the corpus size (number of retrieval units in the corpus). Then, the retriever’s task becomes much easier with more complete information.
@@
    Long Retriever: The long retriever will identify coarse relevant information for the given query by searching through all the long retrieval units in the corpus. Only the top 4 to 8 retrieval units (without re-ranking) are used for the next step.
@@
    Long Reader: The long reader will further extract answers from the concatenation of retrievals, which is normally around 30K tokens. We simply prompt an existing long-context LM (like Gemini or GPT4) with the question to produce the answers in a zero-shot fashion.
These three novel designs significantly boost the overall performance of RAG on open-domain question-answering tasks like NQ Kwiatkowski et al. (2019) and HotpotQA Yang et al. (2018). LongRAG has several advantages: 1) It does not require additional re-rankers and the best results can be attained by only considering the top 4-8 retrieved units. 2) The long retrieval unit amalgamates comprehensive information from related documents, which can be used directly to answer multi-hop questions without iterative retrieval.
In our experiments, we adopt off-the-shelf retrievers like BGE Xiao et al. (2023) and readers like Gemini-1.5-Pro (Reid et al., 2024) or GPT-4o OpenAI (2024) without any tuning on NQ or HotpotQA. In our experiments, we reduce the NQ corpus size from 22M to 600K document units, which improves the answer recall@1 from 52% (DPR) to 71%. Similarly, we reduce the HotpotQA corpus size from 5M to 500K, which improves the recall@2 from 47% (DPR) to 72%. The improvement in retriever can significantly benefit the reader model. By exploiting the long-context understanding ability of GPT-4o, LongRAG can achieve an EM of 62% on NQ and 64% on HotpotQA. These results could be comparable to the strongest fully trained RAG models like Atlas Izacard et al. (2022) and MDR Xiong et al. (2020b).
We perform ablation studies in subsection 4.4 to prove why longer retrieval units are necessary. Given a budget of 40K recall tokens, with ‘short retriever units’, we can increase the number of recalled units to reach a marvelously high recall score (91% for recall@200). However, the end performance dips significantly due to the huge amount of ‘hard negatives’, which confuses the reader. With ‘long retriever units’, we observe an entirely different trend. As we recall more units (from 1 to 8 units), both the recall and end performance will increase or plateau. The impact of ‘hard negative’ is much less severe in LongRAG. It shows that LongRAG can better exploit the advances in the long-context LLMs (reader). As the long-context methods evolve, the performance of LongRAG will continue to improve. Therefore, we believe the modern RAG systems should re-consider the granularity of their retrieval units to exploit the advantages of the current long-context LLMs.
Meanwhile, there is still room for improvement in our framework, particularly the need for stronger long embedding models, as shown in Table 3. Additionally, more general methods to formulate long retrieval units beyond hyperlinks will be helpful.
@@ 3.1 Long Retriever
We approximate it by maximizing the scores of all chunks $g^{\prime}$ within the retrieval unit $g$, akin to the MaxP design in Dai and Callan (2019). We consider different levels of granularity, including passage level, document level, and the complete grouped document. The empirical study about this settings is in Table 3. With this similarity score setup, we will retrieve the top $k$ retrieval units closest to the given query. For efficient retrieval, we precompute the embedding of each retrieval unit $g^{\prime}$ and predict the exact inner product search index in FAISS (Johnson et al., 2019).
@@ 4.2 Retrieval Performance
BGE-Large 512-tokens chunk 71.7%
E5-Mistral-7B 4000-tokens chunk 54.2%
E5-Mistral-7B entire grouped retrieval unit 23.4%
@@ 4.3 Full QA Performance
Figure 5: This table shows the end performance of different granularity of retrieval units. The green curve indicates the ‘large unit’ adopted in LongRAG. The number above the curve indicates the recall. It can be seen that end-performance does not increase monotonically with recall score.
4.4 Ablation Studies
We perform several in-depth ablation to understand what are the important factors in our LongRAG system including "unit size" and "reader variant".
@@ 4.4 Ablation Studies
Recall vs. EM
In Figure 5, we compare the relationship between retrieval recall and end performance across varying context lengths for different retrieval unit selections. We observe that using fewer retrieval units in the reader with longer retrieval units design reduces the introduction of distractors or hard negatives under a given length budget. Consequently, the end performance does not increase monotonically with the recall score. In the future, with advancements in long embedding models and improved retrieval recall for long retrieval units, we can expect better end performance.
@@
In Figure 6, we compare the performance of six different readers: Gemini-1.5-pro, GPT-4-Turbo, GPT-4o, Claude-3-Opus, Claude-3.5-Sonnet and DeepSeek-V2-Chat. The results indicate that GPT-4o achieves the highest exact match score on the 200 test questions of the NQ dataset among the three models. This suggests that GPT-4o is the most effective in the role of a long reader in the LongRAG framework. The enhanced performance of GPT-4o can be attributed to its superior ability to process and comprehend lengthy contexts, ensuring that crucial information is accurately extracted. Therefore, we mainly report the GPT-4o results in our main table. Besides, Gemini-1.5-pro, GPT-4-Turbo, Claude-3-Opus, and Claude-3.5-Sonnet could achieve very similar results. These state-of-the-art black box LLMs are also effective readers within the LongRAG framework. Deepseek-V2-Chat is one of the best open-source LLMs, but its performance degrades significantly compared to the previous five black-box LLMs. The above experiments demonstrate that our current framework depends on the long-context understanding ability of LLMs, and we still have a long way to go in harnessing open-source LLMs within our framework.